# Emergence of Fluctuation Relations in UNO

### v1

@@ Appendix C Crazy Eights
Figure 7: The associated plots for the game Crazy Eights, of which UNO is based.
UNO is derived from a simpler game played with the standard playing card deck called the Crazy Eights. In the most basic version of the game, the eights act as the wildcards and the suites play the role of the colours. More homebrew rules can be added that make the game more similar to UNO. For completeness we create similar plots for the basic Crazy Eights game, which can be seen in Fig. 7. It can be seen that the plots are essentially the same, especially with UNO without the plus-cards (see Fig. 3(b)).
Appendix D A different kind of bath
Figure 8: The Crooks’ plot for an UNO game with other players playing smartly. Compare with Fig. 3(a).
In the main text, the other players other than Alice always play randomly without any strategy. This begs the question of whether this randomness of the other players ‘adds’ to the stochasticity to the bath and affects our result. The short answer to this question is: no. Fig 8 shows the Crooks’ plot obtained with the same method as in the main text but with all the other players using Algorithm 1 instead of playing randomly. As can be seen, there is little to no noticeable difference with Fig. 3(a).

### v2

@@ Appendix D Crazy Eights
Figure 10: The associated plots for the game Crazy Eights, on which UNO is based. Each of the data for these variants was obtained from two million runs.
UNO is derived from a simpler game played with the standard playing card deck called the Crazy Eights. In the most basic version of the game, the eight cards act as the wildcards and the suites play the role of the colors. Additional homebrew rules can be added into the game that makes it closer to UNO. For completeness we create similar plots for the basic Crazy Eights game, which can be seen in Fig. 10. It can be seen that the plots are very similar to UNO, especially when the plus-cards are removed [see Fig. 3(b)].